OBSERVATION AND ROTATIONAL ANALYSIS OF BANDS IN THE C′′ ^{5}Π_{ui}–A′ ^{5}Σ_{g}^{+} SYSTEM OF N_{2}

\underline{O. PIRALI}^{a} and D. W. TOKARYK, Department of Physics and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, Canada.

The singlet and triplet electronic states of N_{2} have been extensively studied, but the quintet manifold has been relatively unexplored. We developed an experimental set-up in order to record the high resolution optogalvanic absorption spectrum of the C′′ ^{5}Π_{ui}–A′ ^{5}Σ_{g}^{+} (Herman InfraRed) system of N_{2} excited in a discharge through a supersonic jet. This very sensitive technique has allowed us to observe of all the spin orbit components of the C′′ ^{5}Π_{ui} state for the first time. A partial analysis of the spectra has been published, which provides the separation between the spin-orbit components F_{1} and F_{2} of the C′′ ^{5}Π_{ui} state.^{b} The experimental value has been confirmed using semi-empirical calculations^{c} based on the spin-orbit parameters of previously observed states of N_{2} (see next presentation). In this talk, we will present the experimental set-up used for this study, the recent complete analysis of our spectra including all five spin-orbit components of the upper state, and a discussion about the values of the fitted parameters.

apresent address: FOM Rijnhuizen, 3439MN Nieuwegein, The Netherlands

bO. Pirali and D. W. Tokaryk, J. Chem. Phys. 124, 081102 (2006)

cR. W. Field, O. Pirali and D. W. Tokaryk, J. Chem. Phys. 124, 081103 (2006)